# Where Do People Tell Stories Online? Story Detection Across Online Communities

### v2

@@ Abstract
StorySeeker toolkit, including a richly annotated dataset of 502 Reddit posts and comments, a detailed codebook adapted to the social media context, and models to predict storytelling at the document and span level.
@@
We evaluate a range of detection methods using our data, and we identify the distinctive textual features of online storytelling, focusing on storytelling span detection, which we introduce as a new task.
@@
Maria Antoniak Joel Mire Maarten Sap Elliott Ash Andrew Piper Allen Institute for AI Carnegie Mellon University ETH Zürich McGill University
@@ 1 Introduction
Several challenges have hindered the large scale analysis of storytelling online. First, defining storytelling, i.e., what is a story and what is not a story, is a difficult task that the field of narratology has been concerned with for decades (Bal and Van Boheemen, 2009). Second, unlike traditional media where titles, genres, and other paratextual cues can signal its presence, storytelling in online communities can be fluid and co-occur with many other kinds of discourse (such as seeking information, providing advice; Yang et al., 2019a). Third, existing story datasets for training and evaluation of a story detector are not publicly available (Ganti et al., 2022, 2023), contain sparse positive labels (Gordon and Swanson, 2009), are not available in English (dos Santos et al., 2017), or include topical confounders (Prize, 2019). Finally, no storytelling detection work has tackled detecting stories embedded within documents.
@@
Im trying to lose 4kg over 2 months or 0.5kg a week. The caloric intake calculator says i should aim for about 2560 calories per day. It also breaks it down into three categories carbs (gm) protein (gm) and fats (gm), which it says is 333 147 91 respectively. Am i right in assuming that (gm) is grams because ive only ever seen it (g). And if so is this about accurate? If it is it makes it easier for me because all the foods here say on the label how many grams of these 3 macros are in the food. I understand this might be a bit basic but [ive searched around the internet and here but havent found an answer though ive probs missed it]. |
Our codebook allows for very short stories composed of as few as two meaningfully connected events. By annotating these minimal spans of narrativity within larger texts, downstream modelling and application can be more attuned to the ways that stories are intermixed with other discourses. Additionally, we can more precisely analyze which textual features are associated with story spans (vs. a less precise approach that considers textual features across the entire text). |
@@
I’m a grad student for context. [It starts out very general, you will spend a year or two really just catching up on literature since, you know, your advisor has been along for the ride for about 15 years so comparatively you’re an academic infant. So you read a paper which answers questions, but also creates more questions for you. But that’s ok and is par for the course, so you look up more papers which answer those questions, and this is a cycle for your entire career. At some point you will think, "why didn’t they do this?" but this time… this wondrous time, you find there is no further information on the topic. But it’s a good thing you spent the past two 2 years catching up on it, and now hopefully you can try and answer the question yourself. Now you read others papers to get answers, which leads to more questions. Suddenly it’s your own set of questions and answers that drive the field.] I am more typical I think and would say after about 12-18 months I started asking questions that didn’t have really well worked out answers, and that’s when the real work begins! I think this diagram is a perfect explanation of how it goes. |
Although prior work on literary event detection focused solely on realis events Sims et al. (2019), which are asserted to have actually happened, our codebook recognizes future-tense and hypothetical stories in certain cases, as in this case, where an author’s experience informs a story-like sequence of events told in a hypothetical grammatical mode. By accounting for these cases in our codebook, we support a more capacious notion of storytelling, which is not limited to a specific grammatical mode. |
@@
Hey all, guess I’ll be coming here more often from now on. I finally ordered an Xbox. I also bought an HDMI adapter on Amazon. Anything I should know? I heard about the undervolting, so I might try that. |
This example demonstrates that the presence of multiple events does not always imply that there is a story. Stories must contain a sequence of events that are meaningfully related. |
@@
Unlike prior work on story detection, (1) we release a public, detailed codebook that can be used across a variety of data types, (2) we explicitly tie our story definition to the related task of event detection (Sims et al., 2019; Vauth et al., 2021) by annotating event spans,
(3) our codebook handles token-level boundaries between storytelling and non-storytelling spans, and (4) our data includes large and diverse sets of online communities rather than focusing on a single setting.
@@
Prior work has not been able to answer fundamental questions such as where storytelling happens online, which community features lead to more storytelling, or how storytelling is used rhetorically across social contexts —
@@ 2 Related Work
In the field of narrative theory (“narratology”), storytelling has been defined by its emphasis on sequences of events, aspects of change and/or conflict, and embodiment or “feltness” Herman (2009); Bruner (1991); Fludernik (2002). As Herman (2009) writes, “Narrative roots itself in the lived, felt experience of human or human-like agents interacting in an ongoing way with their surrounding environment.” Piper et al. (2021b) propose a minimum schema for capturing storytelling that emphasizes the presence of ten basic elements, including an agent and action.
@@
Operationalizing storytelling schema for NLP is an active area of research (Roos and Reccius, 2021; Piper et al., 2021a; Piper and Bagga, 2022). NLP research on narratives and storytelling has deployed a wide range of definitional dimensions, the most common being sequences of events arranged temporally, causally related events leading to resolutions, and entities or characters, while a smaller number include a rhetorical purpose for the text and world building (Ceran et al., 2012; Yao and Huang, 2018; Eisenberg and Finlayson, 2017; Castricato et al., 2021; Alzahrani et al., 2016). See Appendix A.6 for an enumeration of story definition features used in prior work.
Prior work in story annotation has mostly focused on specific discourse domains such as healthcare (Ganti et al., 2022), argumentation (Falk and Lapesa, 2022, 2023), book publishing (Piper and Bagga, 2022), or blogs (Eisenberg and Finlayson, 2017; dos Santos et al., 2017). To our knowledge, all prior work on story detection has focused on passage- or document-level annotations. Additionally, much past annotation work is not publicly available (Gordon and Swanson, 2009; Ceran et al., 2012; Ganti et al., 2023), due to data agreements, sensitive content, and other constraints. We expand on these works by widening our view to many communities and topics, by using a span-based approach, and by publicly releasing our annotations.
@@
Most prior work on automatic story detection has focused on using feature-based classification approaches, relying on features like n-grams, POS tags, and coreference chain length (Ceran et al., 2012; Gordon and Swanson, 2009; Yao and Huang, 2018; Eisenberg and Finlayson, 2017; Piper et al., 2021a). See Appendix A.6 for an enumeration of story features used for prediction in prior work. These studies either had an explicit goal of using interpretable methods or were conducted prior to the arrival of large language models (LLMs). Two newer works have attempted narrative detection using LLMs (Ganti et al., 2022, 2023). Both studies hand-annotate a small set of texts from online healthcare communities with binary story labels, and both find that fine-tuned BERT-based models perform better than classical models at the document-level story prediction task.
@@ 3 Designing a Story Codebook
3.1 Codebook Definitions
Our definition must work across diverse online communities while recognizing features that make storytelling in these contexts potentially unique from storytelling in literary or visual contexts.
Our interdisciplinary team worked iteratively to code data, compare labels, and design a codebook containing a story definition that fit our context (diverse online communities) and research goal (measuring storytelling). Drawing from prior work, we focused our story definition on agent-centered events, leading to a two-step annotation process: first, the annotation of event spans, and second, the annotation of story spans. We provide our full codebook instructions in Appendix A.7.
Our strong emphasis on events while annotating story spans is novel and adds consistency to the annotation of a diverse set of texts. It also allows us to build on prior work on event span annotation (Sims et al., 2019) by reconsidering this task within the context of storytelling, leading us to a more flexible definition of events than prior work.
Story Definition
A story describes a sequence of events involving one or more people. Notably, as long as a text meets these requirements, we annotate it as a story, regardless of whether it is as short as one sentence or as long as the entire post. Storytelling is thus not limited to only posts that are entirely focused on storytelling. Unlike some prior work, we do not include world-building or setting in our definition, as such descriptions are rare in our data. When selecting story spans, we do not include the post title, introductory text about the subreddit, or explanations and discussions external to the story, but we do include non-event text that sets the stage, summarizes the story, or ends with a lesson learned.
Event Definition
Our events definition draws heavily from the event definition in Sims et al. (2019), summarized as an event is a singular occurrence at a particular place and time. We modify this definition in three ways: (1) We do not require verbs to be in the past tense, (2) we sometimes allow hypothetical verbs to be labeled as events, and (3) we sometimes allow verbs with negation to be labeled as events. These changes reflect both the different norms in online communities (e.g., the frequency of using present tense to tell stories), and the story-detection intention underlying our event labeling (e.g., the importance of some negative events in a story sequence). Some changes, like the increased attention to stative verbs, reflect choices also made by Vauth et al. (2021), a follow-up work on literary event detection.
@@ 4 Creating a Multi-Community Corpus
Sampling across the 500 most frequent subreddits in the dataset, we follow an open-coding approach to categorize the subreddits into 33 categories (e.g. professional advice). We use these categories to remove sensitive (e.g., r/confessions), toxic (e.g., r/pettyrevenge), explicit (e.g., r/sex), and non-English (e.g., r/mexico) subreddits from the human annotation tasks, and we use these categories to structure our analyses (§6). We show the full set of categories and their member subreddits in Appendix A.2. We sample a balanced set of five texts from each of the filtered subreddits, downsampling gaming222The gaming category has twice the number of subreddits of the next most frequent category, hobbies (100 versus 49). and requiring that each text contain at least 100 tokens and no more than 500 tokens.
@@
Our annotation process included highlighting both story and event spans, as we found that first identifying events was crucial in making the story span decision. After several rounds of codebook construction, two of the authors independently annotated the target data using Prodigy.333https://prodi.gy/ Annotation was challenging, as texts can contain many events, and we included frequent rounds of group deliberation for difficult examples as well as a final round of discussion to arrive at a single consensus label for each document.
@@
post |
story |
137 |
64% |
119 (mean) |
| non-story |
78 |
36% |
comment |
story |
98 |
34% |
92 (mean) |
| non-story |
189 |
66% |
@@ 5 Developing a Story Detection Model
| Document Classification | Story () | | Not Story () | ||||
@@
Model Training
Following prior work, we evaluate a series of models for our story detection task, including a SVMs baseline using TF-IDF features, a fine-tuned RoBERTa444We used the HuggingFace library with the roberta-base model for both document and sequence prediction. model (Liu et al., 2019), and zero-shot and few-shot GPT-4 prompting (OpenAI, 2023). For few-shot tests with OpenAI models, we prompt with two positive and two negative examples in the prompt, using model versions GPT-4-0314 and GPT-3.5-turbo-0613. We divide our expert-annotated data into a training/prompting set of 301 texts, a validation set of 100 texts, and a test set of 101 texts. We rely on the consensus labels, i.e., if the annotators agree after discussion that a text contains a story span, then we use this as either (i) a positive instance of storytelling or (ii) an indicator of whether to include the union of the annotator’s spans for token prediction.
@@
We find the best story-detection performance from the GPT-4 zero-shot and few-shot prompts (see Table 4), but the fine-tuned RoBERTa classifier is comparable. GPT-4 performed better than GPT-3.5 and chain-of-thought prompting (Camburu et al., 2018) did not yield consistent improvements. Given our small expert set, these results are averaged over bootstrapped samples of the test set, and we show both the means and standard deviations. We examine model errors in Appendix A.1.
@@
Figure 1 shows the results across all story-annotated datasets to which we had access.
We find that our training data generalizes to these non-Reddit datasets reasonably well, with lowest performance for the set of Brazilian blog posts (Ceran et al., 2012), a dataset that we automatically translated from Portuguese to English for this experiment.
The PiperBagga datasets (Piper and Bagga, 2022) include both a full dataset, with a large number of automatic genre-based labels, and a smaller dataset containing 394 hand-annotated texts mostly from literary and non-social media sources.
The other datasets do not include span annotations and so we cannot draw comparisons about story spans.

### v3

@@ Abstract
StorySeeker toolkit, including a richly annotated dataset of 502 Reddit posts and comments, a detailed codebook adapted to the social media context, and models to predict storytelling at the document and span levels.
@@
We evaluate a range of detection methods using our data, and we identify the distinctive textual features of online storytelling, focusing on storytelling spans.
@@
Maria Antoniak♣ Joel Mire♢ Maarten Sap♢♣ Elliott Ash♠ Andrew Piper♡ ♣Allen Institute for AI ♢Carnegie Mellon University ♠ETH Zürich ♡McGill University
@@ 1 Introduction
Several challenges have hindered the large scale analysis of storytelling online. First, defining storytelling, i.e., what is a story and what is not a story, is a difficult task that the field of narratology has been concerned with for decades (Bal and Van Boheemen, 2009). Second, unlike traditional media where titles, genres, and other paratextual cues can signal its presence, storytelling in online communities can be fluid and co-occur with many other kinds of discourse (such as seeking information, providing advice; Yang et al., 2019a). Third, existing story datasets are not appropriate for training and evaluating a cross-community story detector; some datasets are not publicly available (Ganti et al., 2022, 2023), others contain sparse positive labels (Gordon and Swanson, 2009), some are not available in English (dos Santos et al., 2017), some include topical confounders (Prize, 2019), and some are largely related to book-based narratives (Piper and Bagga, 2022). Finally, no storytelling detection work has tackled detecting stories embedded within documents.
@@
| Im trying to lose 4kg over 2 months or 0.5kg a week. The caloric intake calculator says i should aim for about 2560 calories per day. It also breaks it down into three categories carbs (gm) protein (gm) and fats (gm), which it says is 333 147 91 respectively. Am i right in assuming that (gm) is grams because ive only ever seen it (g). And if so is this about accurate? If it is it makes it easier for me because all the foods here say on the label how many grams of these 3 macros are in the food. I understand this might be a bit basic but [ive searched around the internet and here but havent found an answer though ive probs missed it]. | Our codebook allows for very short stories composed of as few as two meaningfully connected events. By annotating these minimal spans of narrativity within larger texts, downstream modelling and application can be more attuned to the ways that stories are intermixed with other discourses. Additionally, we can more precisely analyze which textual features are associated with story spans (vs. a less precise approach that considers textual features across the entire text). |
@@
| I’m a grad student for context. [It starts out very general, you will spend a year or two really just catching up on literature since, you know, your advisor has been along for the ride for about 15 years so comparatively you’re an academic infant. So you read a paper which answers questions, but also creates more questions for you. But that’s ok and is par for the course, so you look up more papers which answer those questions, and this is a cycle for your entire career. At some point you will think, "why didn’t they do this?" but this time… this wondrous time, you find there is no further information on the topic. But it’s a good thing you spent the past two 2 years catching up on it, and now hopefully you can try and answer the question yourself. Now you read others papers to get answers, which leads to more questions. Suddenly it’s your own set of questions and answers that drive the field.] I am more typical I think and would say after about 12-18 months I started asking questions that didn’t have really well worked out answers, and that’s when the real work begins! I think this diagram is a perfect explanation of how it goes. | Although prior work on literary event detection focused solely on realis events Sims et al. (2019), which are asserted to have actually happened, our codebook recognizes future-tense and hypothetical stories in certain cases, as in this case, where an author’s experience informs a story-like sequence of events told in a hypothetical grammatical mode. By accounting for these cases in our codebook, we support a more capacious notion of storytelling, which is not limited to a specific grammatical mode. |
@@
| Hey all, guess I’ll be coming here more often from now on. I finally ordered an Xbox. I also bought an HDMI adapter on Amazon. Anything I should know? I heard about the undervolting, so I might try that. | This example demonstrates that the presence of multiple events does not always imply that there is a story. Stories must contain a sequence of events that are meaningfully related. |
@@
In doing so, (1) we release a public, detailed codebook that can be used across a variety of data types, (2) we explicitly tie our story definition to the related task of event detection (Sims et al., 2019; Vauth et al., 2021) by annotating event spans,
(3) our codebook handles token-level boundaries between storytelling and non-storytelling spans, and (4) our data includes large and diverse sets of online communities rather than focusing on a single online setting, furthering our goal of providing a generalizable story annotation framework.
@@
Prior work has not been able to answer important questions such as where storytelling happens online, which community features lead to more storytelling, or how storytelling is used rhetorically across social contexts —
@@ 2 Related Work
In the field of narrative theory (“narratology”), storytelling has been defined by its emphasis on sequences of events, aspects of change and/or conflict, and embodiment or “feltness” Herman (2009); Bruner (1991); Fludernik (2002). As Herman (2009) writes, “Narrative roots itself in the lived, felt experience of human or human-like agents interacting in an ongoing way with their surrounding environment.” Piper et al. (2021b) propose a minimum schema for capturing storytelling that emphasizes the presence of ten basic elements, including an agent, action, and location in time and space.
@@
Operationalizing storytelling schemas for NLP is an active area of research (Roos and Reccius, 2021; Piper et al., 2021a; Piper and Bagga, 2022). NLP research on narratives and storytelling has deployed a wide range of definitional dimensions, the most common being sequences of events arranged temporally, causally related events leading to resolutions, and the presence of entities or characters, while a smaller number include a rhetorical purpose for the text and world building (Ceran et al., 2012; Yao and Huang, 2018; Eisenberg and Finlayson, 2017; Castricato et al., 2021; Alzahrani et al., 2016). See A.6 for an enumeration of story definition features used in prior work.
Prior work in story annotation has mostly focused on specific discourse domains such as healthcare (Ganti et al., 2022), argumentation (Falk and Lapesa, 2022, 2023), book publishing (Piper and Bagga, 2022), bereavement (Doyle et al., 2024), or blogs (dos Santos et al., 2017; Gordon and Swanson, 2009). To our knowledge, all prior work on story detection has focused on passage- or document-level annotations, except for a set of preliminary annotation guidelines for embedded narratives without an associated dataset by Eisenberg and Finlayson (2021) and a domain-specific study by Falk and Lapesa (2023). Much past annotation work is not publicly available (Ceran et al., 2012; Ganti et al., 2023), due to data agreements, sensitive content, and other constraints.
We expand on these works by widening our view to many communities and topics, by using a span-based approach, and by publicly releasing our annotations. Our annotation guidelines presented in §3 most closely resemble the guidelines provided by Eisenberg and Finlayson (2017) and Eisenberg and Finlayson (2021), which rely on events and characters, and draw most inspiration from the narrative annotation guidelines provided by Piper et al. (2021a) and the event annotation guidelines provided by Sims et al. (2019).
@@
Most prior work on automatic story detection has focused on using feature-based classification approaches, relying on features like n-grams, POS tags, and coreference chain length (Ceran et al., 2012; Gordon and Swanson, 2009; Yao and Huang, 2018; Eisenberg and Finlayson, 2017; Piper et al., 2021a). See A.6 for an enumeration of story features used for prediction in prior work. These studies either had an explicit goal of using interpretable methods or were conducted prior to the arrival of large language models (LLMs). Two newer works have attempted narrative detection using LLMs (Ganti et al., 2022, 2023). Both studies hand-annotate a small set of texts from online healthcare communities with binary story labels, and both find that fine-tuned BERT-based models perform better than classical models at the document-level story prediction task.
@@ 3 Designing a Story Codebook
Our guidelines must work across diverse online communities while recognizing features that make storytelling in these contexts potentially unique from storytelling in literary or visual contexts.
Our interdisciplinary team worked iteratively to code data, compare labels, and design a codebook containing a story description that fit our context (diverse online communities) and research goal (measuring storytelling). Drawing from prior work, we focused our story identification guidelines on agent-centered events, leading to a two-step annotation process: first, the annotation of event spans, and second, the annotation of story spans. We provide our full codebook instructions in A.7.
Our strong emphasis on events while annotating story spans is novel and adds consistency to the annotation of a diverse set of texts. It also allows us to build on prior work on event span annotation (Sims et al., 2019) by reconsidering this task within the context of storytelling, leading us to a more flexible description of events than prior work.
Importantly, our codebook should be considered not as presenting complete or final definitions of either stories or events but rather providing actionable advice that captures features useful for the consistent annotation of these phenomena.
Story annotation guidelines
During our story annotation process, we look for texts containing a sequence of events involving one or more people. Notably, as long as a text meets these requirements, we annotate it as a story, regardless of whether it is as short as one sentence or as long as the entire post. Storytelling is thus not limited to only posts that are entirely focused on storytelling. Unlike some prior work, we do not include world-building or setting as features in our guidelines, as such descriptions are rare in our data, and we also do not include the presence of a narrator, as all Reddit posts and comments by default have their authors as narrators. When selecting story spans, we do not include the post title, introductory text about the subreddit, or explanations and discussions external to the story, but we do include non-event text that sets the stage, summarizes the story, or ends with a lesson learned.
Event annotation guidelines
Our events guidelines draw heavily from the guidelines in Sims et al. (2019), summarized as an event is a singular occurrence at a particular place and time. We modify this guideline in three ways: (1) We do not require verbs to be in the past tense, (2) we sometimes allow hypothetical verbs to be labeled as events, and (3) we sometimes allow verbs with negation to be labeled as events. These changes reflect both the different norms in online communities (e.g., the frequency of using present tense to tell stories), and the story-detection intention underlying our event labeling (e.g., the importance of some negative events in a story sequence). Some changes, like the increased attention to stative verbs, reflect choices also made by Vauth et al. (2021), a follow-up work on literary event detection.
@@ 4 Creating a Multi-Community Corpus
Sampling across the 500 most frequent subreddits in the dataset, we follow an open-coding approach to categorize the subreddits into 33 categories (e.g. professional advice). We use these categories to remove sensitive (e.g., r/confessions), toxic (e.g., r/pettyrevenge), explicit (e.g., r/sex), and non-English (e.g., r/mexico) subreddits from the human annotation tasks, and we use these categories to structure our analyses (§6). We show the full set of categories and their member subreddits in A.2. We sample a balanced set of five texts from each of the filtered subreddits, downsampling gaming222The gaming category has twice the number of subreddits of the next most frequent category, hobbies (100 versus 49). and requiring that each text contains at least 100 tokens and no more than 500 tokens.
@@
Our annotation process included highlighting both story and event spans, as we found that first identifying events was crucial in making the story span decision. After several rounds of codebook construction, two of the authors independently annotated the target data using Prodigy; each annotator annotated the full set of texts.333https://prodi.gy/ Annotation was challenging, as texts can contain many events, and we included frequent rounds of group deliberation for difficult examples as well as a final round of discussion to arrive at a single consensus label for each document.
@@
| post | story | 137 | 64% | 119 (mean) |
| non-story | 78 | 36% | – | |
| comment | story | 98 | 34% | 92 (mean) |
| non-story | 189 | 66% | – | |
@@ 5 Developing a Story Detection Model
| Document Classification | Story | | Not Story | ||||
@@
Prediction methods
To predict the presence of storytelling, we evaluate a variety of prediction models and methods to cover reasonable baselines that vary in cost, GPU usage, and need for labeled data. These methods include a SVMs baseline using TF-IDF features, a fine-tuned RoBERTa444We used the Hugging Face library with the roberta-base model for both document and sequence prediction, using RobertaForSequenceClassification for document prediction and RobertaForTokenClassification for sequence prediction from the transformers library. model (Liu et al., 2019), and zero-shot and few-shot GPT-4 prompting (OpenAI, 2023).
For the document classification task, we predict the binary presence of storytelling in the document, while for the span detection task, we predict whether each token is part of a story. We divide our expert-annotated data into a training/prompting set of 301 texts, a validation set of 100 texts, and a test set of 101 texts. We rely on the consensus labels, i.e., if the annotators agree after discussion that a text contains a story span, then we use this as either (i) a positive instance of storytelling or (ii) an indicator of whether to include the union of the annotator’s spans for token prediction.
We experimented with different prompts using our validation set, including variations of the below task that included examples, chain-of-thought questions, and guidelines.
Task: A story describes a sequence of events involving one or more people. Does the following text contain a story? Answer yes or no, and then explain your reasoning.
Text: <TEXT>
Answer:
For more details and the full set of prompts, see A.8. For few-shot tests with OpenAI models, we interleave two positive and two negative examples in the prompt, using model versions GPT-4-0314 and GPT-3.5-turbo-0613.
@@
We find the best overall story-detection performance from the finetuned RoBERTa model (Table 4), but the GPT-4 prompts are sometimes comparable. GPT-4 performed better than GPT-3.5 and chain-of-thought prompting (Camburu et al., 2018) did not yield consistent improvements. Given our small expert set, these results are averaged over cross-validation folds, and we show both the means and standard deviations. We examine model errors in A.1.
@@
Figure 1 shows the results across the story-annotated datasets to which we could reasonably compare.
We find that our training data generalizes to these non-Reddit datasets reasonably well, with the lowest performance for the set of Brazilian blog posts (Ceran et al., 2012), a dataset that we automatically translated from Portuguese to English using Google Translate for this experiment.
The PiperBagga datasets (Piper and Bagga, 2022) include both a full dataset, with a large number of automatic genre-based labels, and a smaller dataset containing 394 hand-annotated texts mostly from literary sources.
Storytelling in the PiperBaggas datasets is strongly correlated with text genre, resulting in models that are overfit when finetuned and tested on the same datasets.
We do not include comparisons to the ICWSM 2009 Spinn3r Dataset (Burton et al., 2009; Gordon and Swanson, 2009) because of its very low ratio of stories to non-stories, nor do we include comparisons to the Hewlett Essays (Prize, 2019) because of their strong topical confounders (storytelling is almost perfectly correlated with essay prompt). Other datasets were not accessible for comparison. None of the other datasets include span annotations, and so we cannot draw comparisons about story spans.